# Representing local protein environments with atomistic foundation models

### v2

@@ Abstract
The local structure of a protein strongly impacts its function and interactions with other molecules. Therefore, a concise, informative representation of a local protein environment is essential for modeling and designing proteins and biomolecular interactions. However, these environments’ extensive structural and chemical variability makes them challenging to model, and such representations remain under-explored. In this work, we propose a novel representation for a local protein environment derived from the intermediate features of atomistic foundation models (AFMs). We demonstrate that this embedding effectively captures both local structure (e.g., secondary motifs), and chemical features (e.g., amino-acid identity and protonation state). We further show that the AFM-derived representation space exhibits meaningful structure, enabling the construction of data-driven priors over the distribution of biomolecular environments. Finally, in the context of biomolecular NMR spectroscopy, we demonstrate that the proposed representations enable a first-of-its-kind physics-informed chemical shift predictor that achieves state-of-the-art accuracy. Our results demonstrate the surprising effectiveness of atomistic foundation models and their emergent representations for protein modeling beyond traditional molecular simulations. We believe this will open new lines of work in constructing effective functional representations for protein environments.
@@ 1 Introduction
Proteins are complex three-dimensional structures composed of hundreds to thousands of atoms, arranged in a specific manner in space. These molecular architectures as well as their dynamic fluctuations are crucial for the function of a protein. Local environments within a protein are highly diverse, due to the variability of the sequence of amino acids and the fold of the amino-acid chain to a 3D object. Computationally representing these inherently heterogeneous, dynamic, and chemically diverse environments is challenging, yet very important: many key properties of proteins critically rely on these local enviroments, for example the ability of a protein to bind a ligand or perform enzymatic activity. Dynamics of a protein is ultimately rooted in the packing in 3D space, and is generally considered crucial for function. In this context, biomolecular nuclear magnetic resonance (NMR) spectroscopy is a powerful technique, allowing to probe protein dynamics and interactions at atomic resolution.
The principal NMR observables are chemical shifts of certain atomic nuclei endowed with non-zero spin (e.g.,1H, 13C, 15N atoms). The chemical shift of an atom is exquisitely sensitive to its local electronic environment and is critically dependent on local structural features, such as the type of amino acid, the dihedral angles between bonds (and, thus, secondary structure), the orientation of side chains, hydrogen bonding, and solvent exposure. These parameters, averaged over time, impact the chemical shift in a complex manner, and while the chemical shift is exquisitely sensitive to them, interpreting chemical shifts in terms of structure is hampered by this complexity.
There is a strong need for accurate computational prediction of chemical shifts, with enormous potential in several directions: accurately predicted chemical shifts from a given protein structure largely facilitates the assignment of observed NMR signals to individual atoms; moreover, the inverse direction of determining structures and dynamics from experimentally observed chemical shifts hinges on the ability to predict chemical shifts with high accuracy.
Existing chemical-shift prediction approaches neal2003shiftx ; shen2008sparta ; ucbshift2023 use experimentally determined chemical shifts and experimental 3D structures; these structures are parametrized with hand-crafted descriptors to represent the local environment (e.g. dihedral angles, hydrogen bonds, ring-current shifts). Likewise, predictors of the protonation state of titratable groups equally employ hand-crafted features olsson2011propka3 . These tools face limitations in adaptability and accuracy. This highlights the need for more expressive, effective, and generalizable representations of local protein environments.
In computational chemistry, descriptors like Parrinello-Behler symmetry functions behler2011atom ; behler2007generalized ; behler2016perspective and Smooth Overlap of Atomic Positions (SOAP) are widely used to represent molecular environments jager2018machine . These methods encode atomic interactions and geometry into concise, invariant representations suited for geometric modeling. Modern atomistic machine learning techniques schutt2018schnet ; schutt2021equivariant implicitly learn similar representations within neural network layers, achieving accuracy comparable to density functional theory (DFT) deng2023chgnet in tasks like molecular dynamics (MD) simulations.
Figure 1: Proposed construction of local protein environment descriptors and their use. A focus residue is used as the center of an environment including full amino acids located in the $5$Å range from it. The atoms are embedded by an atomistic foundational model (AFM), producing a descriptor for each atom in the environment influenced by all the other atoms. The embeddings corresponding to the principal atoms of the focus residue form a canonical descriptor of the environment and may comprise zeroth and higher-order irreducible representation components. These descriptors are then used by downstream models predicting atom, residue, or environment properties.
Recent advances have shifted neural-network interatomic potentials from element-specific models to universal atomistic foundation models [^1] We use “atomistic foundation models” to denote transferable force fields and universal neural-network potentials. Analogous to foundation models in vision and language, they learn compact representations of atomistic systems that can be reused for a variety of downstream tasks as we demonstrate here. (AFMs) trained on large datasets of molecules with DFT-simulated energies. ANI-2x pioneered this transition, covering biologically relevant elements (H, C, N, O, F, S, Cl) with DFT-level accuracy across millions of organic molecules smith2018less . Three recent model families – Multi Atomic Cluster Expansion (MACE) batatia2022mace , OrbNet qiao2020orbnet , and AIMNet zubatyuk2019accurate – have further advanced the field: MACE combines equivariant message passing with high-order many-body terms for materials science applications; OrbNet enhances graph neural networks with semi-empirical orbital features, reducing data requirements for DFT-level accuracy by an order of magnitude; and AIMNet-2 extends ANI with learned atomic embeddings and multitask training (energies, charges, spin), improving transferability to charged and open-shell systems.
In this study, atomistic foundation models are used to describe local protein structure. While typically applied to molecular simulations kapil2024first ; lahey2020benchmarking , our focus is different: analyzing AFM internal representations to reveal the biochemical and biophysical information they encode.
Contributions and Summary of Results
In what follows, we outline our major contributions and provide a summary of results.
@@
Canonical local environments and AFM representations. Protein environments vary widely, making their direct comparison challenging. To make residue-level representations comparable, we introduce canonical environments: regions centered on a focus residue and containing every amino acid whose atom coordinates lie within $5$Å (Hausdorff distance) of a focus residue. We then construct transferable environment representations from the AFM embeddings of the backbone atoms and their bonded hydrogens – N, CA, CB, C, O, H, HA – whose roles are conserved across residues (see Fig. 1). To our knowledge, this is the first use of AFMs to characterize local protein structure.
@@
AFM representations effectively capture local protein structure and chemistry. By analyzing over $165$K protein environments extracted from $1327$ non-redundant chains, we demonstrate that their embeddings reveal secondary-structure features and amino-acid identities (see Fig. A.8). In addition, we train lightweight task networks on these embeddings to predict protonation states obtained from a Poisson–Boltzmann solver reis2020pypka , and find that they achieve higher accuracy than pKa-ANI gokcan2022prediction and propKa olsson2011propka3 (see Table B.3).
@@
AFM representations allow computation of similarity metrics and calculating likelihoods. We show that AFM embeddings have a meaningful latent space. Specifically, distances between environment embeddings reflect both geometric similarity and chemical context, enabling intuitive clustering of related local motifs. We construct likelihood and conditional likelihoods in the AFM embedding space and show that they effectively capture the distribution of biomolecular environments (see Fig. A.95). We show that the defined likelihoods capture subtle structural deformations in protein structure and allow capturing distributions of shifts, making them valuable for structural quality assessment, uncertainty estimation, anomaly detection, dataset curation, and potentially in guiding Boltzmann generators noe2019boltzmann . Furthermore, leveraging the AFM embedding space, we introduce a similarity metric to compare protein environments (see Appendix G.1).
Figure 2: Synthetic example showing the influence of a phenylalanine sidechain aromatic ring on surrounding chemical shifts. (A) The magnitude of change of backbone CA chemical shifts over different ring orientations as predicted using the proposed MACE-based predictor (left) and UCB Shift X (right). A $7$Å sphere indicates the radius from the ring center at which the influence of the ring current is expected to become negligible. Note that UCB Shift X predicts much longer-range, albeit small, ring influence extending beyond $20$Å. (B) Locations of three nearby CA atoms; and (C) their predicted chemical shifts plotted vs. the ring orientation. Note the smooth $180^{\circ}$-periodic behavior of the Mace shift prediction and the decay of the effect scale with the distance from the ring.
@@
AFMs enable physics-grounded accurate chemical shift predictor. In the context of biomolecular NMR, we leverage AFM embeddings to predict chemical shifts in proteins. We demonstrate that a lightweight downstream network receiving these embeddings as the inputs outperforms the current state-of-the-art chemical shift predictor, UCBShift2-X, on backbone and side-chain atoms (see Fig. 4). We conduct a case study examining the ring current effects produced by aromatic side chains on chemical shifts of neighboring nuclei, and demonstrate that our predictor follows physically consistent trends, whereas UCBShift2-X shows unphysical behavior (see Fig. 2). We also pair the predictor with AFM-derived likelihoods to assign confidence scores, yielding a reliable uncertainty measure around the predicted shifts (see Fig. 5).
@@
Interpreting AFM representations. We probe the content of AFM embeddings in two controlled settings. Firstly, we systematically modify side chain orientations and track how the embeddings respond (see Fig. 2 and A.7). Secondly, we follow an unfolding simulation in which a seven-residue $\alpha$-helix extends into a strand (see Fig. A.10A.6). In both cases, principal-component analysis uncovers clear latent directions that mirror the underlying structural changes. We then ask whether an environment descriptor alone suffices to reconstruct the local structure. By optimizing AlphaFold3 abramson2024accurate structures with only this descriptor, we obtain alternative conformations that reproduce the backbone arrangement and many side-chain orientations, although fine-grained details remain imperfect maddipatla2025inverseproblemsexperimentguidedalphafold ; maddipatla2024generativemodelingproteinensembles ; levy2025solvinginverseproblemsprotein . This partial recovery suggests that AFM embeddings encode most—but not all—of the information required to rebuild a protein environment (see Appendix. I.1).
@@
Benchmarking different AFMs. We benchmark three model families—MACE & EGRET, OrbNet, and AIMNet—on tasks defined on protein environment that include secondary-structure assignment, amino-acid identification, acid dissociation constant (pKa) and chemical-shift prediction (see Tables B.2B.1B.4). MACE-based embeddings perform best on every task except pKa prediction, where AIMNet exhibits superior performance (see Table B.3).
These findings show that AFM embeddings compactly and effectively describe complex biomolecular environments, we believe this would inspire future studies of local protein structure with AFMs. The rest of the paper is organized as follows: In Section 2 we provide a brief introduction to AFMs and dedicate Section 3 to the construction of the proposed local environment descriptors. In Section 4, we demonstrate that the proposed representations capture local protein structure and chemistry. In Section 5, we discuss the evaluation of likelihood and similarity of protein environments. Section 6 is dedicated to the prediction of NMR chemical shifts from AFM representations. In Section 7, we discuss the interpretation of the proposed representations. Section 8 briefly overviews related work and Section 9 discusses limitations and future research directions. We encourage the reader to have a detailed look at the main and supplementary figures visualizing the main findings of the paper.
2 Background on Atomistic Foundation Models
Force field models approximate a system’s potential energy from atomic positions and types. The gradient of this energy with respect to atomic positions defines the forces acting on each atom. Traditional force fields were modelled by empirically computing interatomic potentials or using DFTs to analytically determine interatomic forces finnis2003interatomic . More recently, AFMs, have been developed to approximate the energy function and force fields from computational quantum mechanical data. In this work, we use the following AFMs (detailed summary in Table B.5 in Appendix):
MACE batatia2022mace ; kovacs2023mace is based on the body-ordered Atomic Cluster Expansion (ACE) dusson2022atomic with higher-order (many-body) equivariant message passing to capture complex atomic interactions. It constructs atomic energies using symmetry-preserving basis functions and updates them through rotation-equivariant neural interactions. MACE is trained on datasets containing small molecules and tripeptides to predict the energies and forces of DFT relaxation trajectories. Introduced recently, Egret wagen2025egret shares a similar higher-order equivariant message-passing architecture with MACE, but introduces architectural optimizations that improve computational efficiency and scalability.
OrbNet qiao2020orbnet enhances graph neural networks by incorporating quantum-derived features calculated using fast approximate quantum chemistry methods. These features include symmetry-adapted atomic orbital overlaps and elements of the Fock matrix szabo1996modern , which capture important information about the electronic structure. OrbNet is trained on DFT relaxation trajectories and quantum chemistry datasets (containing small molecules and materials) to predict molecular energies and properties.
AIMNet zubatyuk2019accurate uses message passing over atomic environment descriptors and iterative updates inspired by charge equilibration schemes. Through these updates, AIMNet captures both short- and long-range interactions and builds a joint, information-rich representation of each atom in its molecular context. The model is trained on diverse quantum mechanical datasets to predict molecular energies, forces, and other properties across diverse chemical systems.
3 Representing protein local environments with AFMs

### v3

# arXiv:2505.23354v3[q-bio.BM] 27 Feb 2026

@@ Abstract
The local structure of a protein strongly impacts its function and interactions with other molecules. Representing local biomolecular environments remains a key challenge while applying machine learning approaches over protein structures. The structural and chemical variability of these environments makes them challenging to model, and performing representation learning on these objects remains largely under-explored. In this work, we propose representations for local protein environments that leverage intermediate features from machine learning force fields (MLFFs). We extensively benchmark state-of-the-art MLFFs, comparing their performance across latent spaces and downstream tasks, and show that their embeddings capture local structural (e.g., secondary motifs) and chemical features (e.g., amino acid identity and protonation state), organizing protein environments into a structured manifold. We show that these representations enable zero-shot generalization and transfer across diverse downstream tasks. As a case study, we build a physics-informed, uncertainty-aware chemical shift predictor that achieves state-of-the-art accuracy in biomolecular NMR spectroscopy. Our results establish MLFFs as general-purpose, reusable representation learners for protein modeling, opening new directions in representation learning for structured physical systems. Code and data are available at https://github.com/mb012/MLFF_representation.
@@ 1 Introduction
Proteins [^1] Proteins are polymers of amino acids (also called residues). Residues share a common backbone of four heavy atoms (N, CA, C, O) and differ in their side chains. are complex three-dimensional structures composed of hundreds to thousands of atoms, arranged in a specific manner in space. Local environments within a protein are highly diverse due to variability in the amino-acid sequence and the folding of the chain into a 3D structure. These factors govern a protein’s functional mechanisms such as ligand binding, catalysis, and allostery, and make proteins uniquely challenging compared to small molecules. Learning compact and transferable representations of local protein environments is a central challenge for performing machine learning (ML) over biomolecules: the difficulty lies in jointly encoding the local chemical context, including atomic identities, bonds, and subtle biochemical properties, into a consistent and generalizable representation that can transfer across diverse protein modeling tasks.
Classical approaches employ hand-crafted descriptors that partially capture this information by embedding features such as dihedral angles, hydrogen bonds, or electrostatic terms, which often limits generalization across proteins and tasks. In computational chemistry, descriptors like Parrinello-Behler symmetry functions (Behler, 2011; Behler & Parrinello, 2007; Behler, 2016) are widely used to represent molecular environments (Jäger et al., 2018). These methods encode atomic interactions and geometry into concise, invariant representations suited for geometric modeling. Modern atomistic ML techniques (Schütt et al., 2018; 2021) implicitly learn similar representations with neural networks, achieving accuracy comparable to density functional theory (DFT) (Deng et al., 2023) and are used for molecular dynamics (MD) simulations. Recent advances have shifted neural-network interatomic potentials from element-specific models to universal machine learned force-fields (MLFFs) trained on large datasets of molecules with DFT-simulated energies. Modern MLFFs train large neural networks to achieve DFT-level accuracy on millions of calculations, with representative families including AIMNet (Smith et al., 2018), MACE (Batatia et al., 2022; Kovács et al., 2025), and OrbNet (Qiao et al., 2020).
Figure 1: Proposed construction of canonical local protein environment descriptors and their use. Machine learning force field (MLFF) models are pre-trained as energy and force regressors on databases of DFT-calculated energies, enabling them to learn zeroth, first, and second order latent representations of interatomic interactions. We embed local protein environments by extracting latent embeddings from pre-trained MLFFs for a focus residue and all atoms within a 55Å radius of the residue. These embeddings are then mapped onto the atoms of the focus residue to construct canonical environment descriptors which can be used in downstream models for transfer learning to predict diverse chemical properties.
Contributions and Summary of Results. The key idea of this work is to repurpose MLFF embeddings from their conventional role in energy and force prediction to representing local chemical environments in proteins. We show that MLFFs learn compact and generalizable representations of atomistic systems that can be reused for a variety of downstream tasks. MLFF embeddings, unlike sequence-based representations (e.g., ESM (Lin et al., 2023)), are trained on quantum-mechanical data and encode physically grounded features such as bond geometry, torsions, and electronic interactions. Because these embeddings are defined per atom, they naturally transfer to unseen proteins. The atomic building blocks remain the same even when sequences or folds differ, enabling generalization to rare chemistries and out-of-distribution conformations. Analogous to foundation models in vision and language, these representations enable zero-shot and transfer learning, support data-driven priors over biomolecular environments, and position MLFFs as foundation models for structural biology. Outlined below are our major contributions and a summary of results. To assist the reader, we provide a succinct summary of the research questions, experimental designs, and key findings in Table LABEL:tab:exp_overview.
@@
Canonical local environments and MLFF representations. Protein environments vary widely, making their direct comparison challenging. To make MLFF representations comparable across residues and proteins, we introduce canonical environments: regions centered on a focus residue and containing every amino acid whose atom coordinates lie within 55Å (Hausdorff distance) of a focus residue. We then construct transferable environment representations from the atom-wise representations of constituent atoms (or a subset thereof). To our knowledge, this is the first use of MLFFs to characterize local protein structure.
@@
MLFF representations effectively capture local protein structure and chemistry. By analyzing over 165165K protein environments extracted from $1048$ non-redundant chains, we demonstrate that MLFF embeddings meaningfully organize biochemical information without any additional training: unsupervised clustering of these embeddings reveal secondary structure features and amino acid identities (see Fig. 2). Furthermore, we use these embeddings for transfer learning by training networks to predict protonation states calculated with a Poisson-Boltzmann solver (Reis et al., 2020). The resulting models based on MLFF embeddings – especially our best AIMNet-feature model – achieve the lowest errors on simulated pKa values, outperforming classical frameworks such as pKa-ANI (Gokcan & Isayev, 2022) and PropKa (Olsson et al., 2011), as well as GCNs built on learned or ESM-derived embeddings (see Table 1).
@@
MLFF representations allow computation of similarity metrics and calculating likelihoods. We show that MLFF embedding space is well-structured. Specifically, distances between environment embeddings reflect both geometric similarity and chemical context, enabling intuitive clustering of related local motifs. We construct likelihood and conditional likelihoods in the MLFF embedding space and show that they effectively capture the distribution of biomolecular environments (see Fig. A.7A.11). We show that the defined likelihoods capture subtle structural deformations in protein structure and allow capturing distribution shifts, making them valuable for structural quality assessment, uncertainty estimation, anomaly detection, dataset curation, and potentially in guiding Boltzmann generators (Noé et al., 2019). Furthermore, leveraging the MLFF embedding space, we introduce a similarity metric to compare protein environments (see Appendix H.2).
@@
MLFFs enable transfer learning of physics-grounded, uncertainty-aware chemical shift predictor. In the context of biomolecular nuclear magnetic resonance (NMR) spectroscopy, we leverage MLFF embeddings and train downstream networks to predict protein chemical shifts, and demonstrate that this approach outperforms the current state-of-the-art predictor UCBShift2-X for both backbone and side-chain nuclei (see Fig. 3). We conduct a case study examining the ring current effects produced by aromatic side chains on chemical shifts of neighboring nuclei, and demonstrate that our predictor follows physically consistent trends, whereas UCBShift2-X shows unphysical behavior (see Fig. 4). We also pair the predictor with MLFF-derived likelihoods to assign confidence scores, yielding a reliable uncertainty measure around the predicted shifts (see Fig. A.7).
@@
Interpreting MLFF representations. We probe the content of MLFF embeddings in two controlled settings. Firstly, we systematically modify side chain orientations and track how the embeddings respond (see Fig. 4 and 6). Secondly, we follow an unfolding simulation in which a seven-residue $\alpha$-helix extends into a strand (see Fig. A.9, A.8, and A.10). In both cases, principal-component analysis uncovers clear latent directions that mirror the underlying structural changes. Finally, to understand the information contained within the MLFF embeddings, we ask the following question: can we invert the MLFF embedding to recover the underlying local protein environment? Casting this question as an inverse problem, we guide (Maddipatla et al., 2025; 2024; Levy et al., 2025) AlphaFold3 (Abramson et al., 2024) and reasonably recover protein conformations underlying a target MLFF embedding. This partial recovery suggests that MLFF embeddings encode necessary information required to invert a local protein conformation (see Appendix. J.2).
@@
Benchmarking different MLFFs. We benchmark three model families – MACE, OrbNet, and AIMNet – on tasks defined on protein environment that include secondary structure assignment, amino acid identification, acid dissociation constant (pKa) and chemical shift prediction (see Tables B.4B.3B.5B.6). MACE-based embeddings perform best on every task except pKa prediction, where AIMNet exhibits superior performance (see Table 1).
Together, these steps recast MLFFs as general‑purpose biochemical feature extractors for proteins, providing canonical representations, principled similarity/likelihood tools, and state‑of‑the‑art downstream performance with calibrated uncertainty. The remainder of the paper is organized as follows. Section 2 introduces MLFFs and reviews closely related work, and Section 3 describes the construction of canonical local environments. Sections 4-6 demonstrate that the proposed representations capture protein structure and chemistry and enable state-of-the-art predictors. Section 7 interprets the learned representations, and Section 8 outlines limitations and future directions. Due to space constraints, the Appendix contains a significant amount of supplementary material, including additional figures, tables, experiments, an extended background, and full experimental details. We encourage the reader to consult the table of contents in the Appendix for a complete overview.
2 Background & Related Work
MLFFs. Machine learning force-fields (MLFFs) are neural networks trained to reproduce quantum-mechanical potential energy surfaces. Given atomic coordinates and elements as inputs, MLFFs apply symmetry-preserving encodings (e.g., local environments, equivariant basis functions), and output per-atom energy contributions whose sum yields the total potential energy. Consequently, forces are obtained as the negative energy gradient, enabling molecular dynamics (MD) simulations with near-quantum accuracy at orders of magnitude lower cost than direct DFT calculations. In this work, we focus on three representative MLFF families with distinct design principles: MACE (Batatia et al., 2022) and Egret (Wagen et al., 2025) combine equivariant message passing with high-order many-body interatomic potentials; OrbNet (Qiao et al., 2020) augments graph neural networks with semi-empirical orbital features to achieve DFT-level accuracy with reduced data requirements; and AIMNet (Zubatyuk et al., 2019) extends ANI with learned atomic embeddings and multitask training (energies, charges, spin), improving transferability to charged and open-shell systems. While these models achieve remarkable accuracy for small molecules, materials, and bipeptides, MLFFs continue to face challenges in capturing long-range interactions in larger biomolecular systems like proteins. Further details on these models and a short background on MD and DFT is provided in Appendix D.
Related work. Recent work has highlighted the versatility of pretrained MLFFs as learned representations for molecular tasks. Shiota et al. (2024) use MACE and M3GNet features to predict chemical shifts in small molecules, while Elijošius et al. (2025) employ MACE embeddings for zero-shot molecular generation via evolutionary search. However these works are limited to small-molecules and do not scale to large, structurally heterogeneous biomolecules like proteins. Similar to our idea, Gokcan & Isayev (2022) extract ionizable residues and use ANI-2x to construct per-residue atomic environment vectors (AEVs), which are then used as input to train a regressor for pKa prediction. However, these AEVs are fixed, hand-crafted descriptors derived from symmetry functions that do not use message passing over the local environment. As a result, they lack the ability to adaptively capture the context-dependent interactions present in protein systems. In contrast, our descriptors encode a richer biochemical context. Moreover, we train a GCN over this many-body embedding space, enabling a context-aware model. As shown in Table 1, our method achieves lower pKa prediction error and generalizes to additional biochemical tasks beyond pKa.
Table 1: Acid Dissociation constant (pKa) prediction accuracy. Reported are mean absolute errors (MAE) with respect to pKa values estimated using the Poisson–Boltzmann solver PypKa (Reis et al., 2020) on protonating / deprotonating amino acids. The table compares (i) classical pKa predictors (Olsson et al., 2011; Gokcan & Isayev, 2022), (ii) a GCN with learned embeddings (“Learned”), (iii) ESM-based features + GCN (Ouyang-Zhang et al., 2025; Hayes et al., 2025; Lin et al., 2023), and (iv) MLFF-based features + GCN (ours). PropKa and pKa-ANI were originally trained on experimental pKa values and are therefore not specifically optimized for the computational reference used here.
The best performing model for each metric is indicated in bold.
ESM-Features + GCN MLFF-Features + GCN (Ours)
Residues # Samples PropKa pKa-ANI Learned ESM3 (Seq) ISM ESM (Struc) ESMFold MACE OrbNet AIMNet Egret
Glutamic acid $12592$ $0.551$ $0.445$ $0.518$ $0.459$ $0.382$ $0.360$ $0.351$ $0.306$ $0.306$ $\mathbf{0.265}$ $0.304$
Aspartic acid $9841$ $0.469$ $0.473$ $0.582$ $0.528$ $0.430$ $0.388$ $0.419$ $0.280$ $0.284$ $\mathbf{0.267}$ $0.272$
Lysine $12009$ $0.393$ $0.401$ $0.652$ $0.359$ $0.304$ $0.295$ $0.278$ $0.320$ $0.282$ $\mathbf{0.270}$ $0.298$
Histidine $1182$ $0.561$ $0.426$ $0.615$ $0.488$ $0.428$ $0.408$ $0.383$ $0.424$ $0.441$ $\mathbf{0.380}$ $0.440$
3 Representing protein local environments with MLFFs